an oopened topped box has a square base (bottom) and a height of 7cm. if the surface area of the five sides of the box is 165 cm squared, find the simensions of the square base.
Answer by venugopalramana(3286)

an oopened topped box has a square base (bottom)....SAY EQUAL TO A
and a height of 7cm. if the surface area of the five sides of the box is 165 cm squared, find the simensions of the square base.
S.A.=BASE AREA +4 SIDES
165=A^2+4*7*A=A^2+28A
A^2+33A-5A-165=0
A(A+33)-5(A+33)=0
(A-5)(A+33)=0
A=5
SQUARE BASE IS 5 AND 5 SIZE